### v1

@@ 4 Discussion
In this paper we have extended the Alcubierre warp drive spacetime to an infinite class of warp drives utilizing the Martel–Poisson charts for Minkowski and AdS spacetimes and their conical defect cousins in three dimensions. The Minkowski spacetime admits Painlevé–Gullstrand coordinates but for AdS spacetime the usual PG charts vanish and the warp drive spacetimes discussed in this paper are the only possibility. Though we have analyzed the warp drives away from the conical singularities at the origin, the fact that spatial hypersurfaces are cones have global consequences which have been captured in the expressions for the expansion and $(-ve)$ energy densities localized around the warp drive boundary.
There are several possibilities for generalizations. We can look for warped spherically symmetric background spacetimes
@@
or generalize the definition of a cone from two to higher dimensions,
@@
where unless $d\tilde{\Omega}_{(D-2)}^{2}$ is the metric of the unit sphere $S^{(D-2)}$, the generalized cone $ds_{C}^{2}$ is singular. For example, MP–charts of various black hole backgrounds would come under its purview. For most of these generalizations, the Ricci scalar of the spatial hypersurface will be non zero thereby opening up new possibilities for the energy density $\rho$. If we focus on the spherically symmetric cases, the expressions involving the extrinsic curvatures would generalize easily from the three dimensional expressions in the previous sections upto some coefficients. We have to symmetrize the expressions in all spatial coordinates except $x$, the direction of travel for the warp drive.
While warp drives are beyond the realm of current experiments, various analog gravity setups imitate the metric Fischer_2003 ; finazzi2012analogue and can be studied experimentally PhysRevLett.105.240401 . It might be possible to generalize these setups to accommodate the Martel–Poisson versions of the warp drive.
Various other properties of the wrap drive are of interest, for example the zero–expansion drives of Natário Nat_rio_2002 , the formation of event horizon and various quantum effects Hiscock_1997 ; Finazzi_2009 ; Barcel__2022 to optimization of various parameters including the form functions and the idea of a warp bubble inside another bubble Broeck_1999 ; Helmerich_2023 . We hope to study some of these aspects in our future works.
@@ Acknowledgements.
The work A.C. is supported by IIT Bhubaneswar Seed Grant SP–103.
@@ Appendix A Different choices for $r_{s}$ and $\gamma_{ij}$
We take $r_{s}=\sqrt{(x-x_{0})^{2}+y^{2}}$ and $\gamma_{ij}=\delta_{ij}+(1-f(r_{s}))\left(\frac{1}{\epsilon^{2}}-1\right)\frac{x^{i}x^{j}}{r^{2}}$. There are two important deviations from the expressions in (3.1). Firstly, the effect of $(1-f(r_{s}))$ term in $\gamma_{ij}$ starts to show only at $\mathcal{O}(v_{\infty}^{2})$, so if we are truncating to $\mathcal{O}(v_{\infty})$, all results will match. Secondly, the spatial curvature tensors, particularly the Ricci scalar ${}^{(2)}R$ is not zero at finite $r$. Its contribution to the energy density $\rho$ starts at order $\mathcal{O}(v_{\infty}^{2})$,
@@
Similarly, the $K_{ij}$ in (20) now picks up contributions from the $\frac{1}{2}\partial_{T}\gamma_{ij}$ term, again modifying expressions from $\mathcal{O}(v_{\infty}^{2})$ onward. In all, this choice complicates things but after background subtraction $\theta$ and $\rho$ are proportional to $f^{\prime}(r_{s})^{\#}$ and $f^{\prime\prime}(r_{s})^{\#}$, therefore localized around the wrap drive bubble. A lot now depends on the choice of $f(r_{s})$, tweaking it might give a $(+ve)$ ${}^{(2)}R$ thereby reducing the overall $(-ve)$ energy density required to operate the drive.
@@
We take $r_{s}=\sqrt{\gamma_{ij}(x-x_{o}(T))^{i}(x-x_{0}(T))^{j}}$ and $\gamma_{ij}=\delta_{ij}+\left(\frac{1}{\epsilon^{2}}-1\right)\frac{x^{i}x^{j}}{r^{2}}$. Here, as we are not changing the spatial metric, there is no contribution from ${}^{(2)}R$, neither $\frac{1}{2}\partial_{T}\gamma_{ij}$ contributes to $K_{ij}$. Extra contributions as compared to results in section 3 are due to the changes in the derivatives of $f(r_{s})$ or equivalently $r_{s}$ w.r.t. $(x,\,y)$ coordinates. Again, the changes are from $\mathcal{O}(v_{\infty}^{2})$ onward and yet again the contributions to $\theta$ and $\rho$ are localized to the boundary of the warp drive.

### v2

@@ 4 Discussion
In this paper we have extended the Alcubierre warp drive spacetime to an infinite class of warp drives utilising the Martel–Poisson charts for Minkowski and AdS spacetimes and their conical defect cousins in three dimensions. The Minkowski spacetime admits Painlevé–Gull strand coordinates but for AdS spacetime the usual PG charts vanish and the warp drive spacetimes discussed in this paper are the only possibility. Though we have analysed the warp drives away from the conical singularities at the origin, the fact that spatial hypersurfaces are cones have global consequences which have been captured in the expressions for the expansion and $(-ve)$ energy densities localised around the warp drive boundary. To the leading order, energy requirements either scale up or down compared to the original Alcubierre warp drive depending upon the choice of the embedding. All the warp drives discussed in the paper violate the null energy condition (NEC).
There are several possibilities for generalisations. We can look for warped spherically symmetric background spacetimes
@@
or generalise the definition of a cone from two to higher dimensions,
@@
where unless $d\tilde{\Omega}_{(D-2)}^{2}$ is the metric of the unit sphere $S^{(D-2)}$, the generalised cone $ds_{C}^{2}$ is singular. For example, MP–charts of various black hole backgrounds would come under its purview. For most of these generalisations, the Ricci scalar of the spatial hypersurface will be non zero thereby opening up new possibilities for the energy density $\rho$. If we focus on the spherically symmetric cases, the expressions involving the extrinsic curvatures would generalise easily from the three dimensional expressions in the previous sections upto some coefficients. We have to symmetrise the expressions in all spatial coordinates except $x$, the direction of travel for the warp drive.
While warp drives are beyond the realm of current experiments, various analog gravity setups imitate the metric Fischer_2003 ; finazzi2012analogue and can be studied experimentally PhysRevLett.105.240401 . It might be possible to generalize these setups to accommodate the Martel–Poisson versions of the warp drive. Various other properties of the wrap drive are of interest, for example the zero–expansion drives of Natário Nat_rio_2002 , the exact analysis of the formation of event horizon and various quantum effects Hiscock_1997 ; Finazzi_2009 ; Barcel__2022 , optimisation of various parameters including the form functions and the idea of a warp bubble inside another bubble Broeck_1999 ; Helmerich_2023 . Another tangent line of investigation would be to explore the AdS/CFT dictionary and its consequences in the warp drive perturbed AdS metrics Maldacena:1997re . We hope to study some of these aspects in our future works.
@@ Acknowledgements.
A.C. thanks Manirujjaman Chowdhury for useful discussions . The work A.C. is supported by IIT Bhubaneswar Seed Grant SP–103.
@@ Appendix A Different choices for $r_{s}$ and $\gamma_{ij}$
We take $r_{s}=\sqrt{(x-x_{0})^{2}+y^{2}}$ and $\gamma_{ij}=\delta_{ij}+(1-f(r_{s}))\left(\frac{1}{\epsilon^{2}}-1\right)\frac{x^{i}x^{j}}{r^{2}}$. There are two important deviations from the expressions in (27). Firstly, the effect of $(1-f(r_{s}))$ term in $\gamma_{ij}$ starts to show only at $\mathcal{O}(v_{\infty}^{2})$, so if we are truncating to $\mathcal{O}(v_{\infty})$, all results will match. Secondly, the spatial curvature tensors, particularly the Ricci scalar ${}^{(2)}R$ is not zero at finite $r$. Its contribution to the energy density $\rho$ starts at order $\mathcal{O}(v_{\infty}^{2})$,
@@
Similarly, the $K_{ij}$ in (22) now picks up contributions from the $\frac{1}{2}\partial_{T}\gamma_{ij}$ term, again modifying expressions from $\mathcal{O}(v_{\infty}^{2})$ onwards. In all, this choice complicates things but after background subtraction, $\theta$ and $\rho$ are proportional to $f^{\prime}(r_{s})^{\#}$ and $f^{\prime\prime}(r_{s})^{\#}$, therefore localised around the wrap drive bubble. A lot now depends on the choice of $f(r_{s})$, tweaking it might give a $(+ve)$ ${}^{(2)}R$ thereby reducing the overall $(-ve)$ energy density required to operate the drive.
@@
We take $r_{s}=\sqrt{\gamma_{ij}(x-x_{o}(T))^{i}(x-x_{0}(T))^{j}}$ and $\gamma_{ij}=\delta_{ij}+\left(\frac{1}{\epsilon^{2}}-1\right)\frac{x^{i}x^{j}}{r^{2}}$. Here, as we are not changing the spatial metric, there is no contribution from ${}^{(2)}R$, neither $\frac{1}{2}\partial_{T}\gamma_{ij}$ contributes to $K_{ij}$. Extra contributions as compared to results in section 3 are due to the changes in the derivatives of $f(r_{s})$ or equivalently $r_{s}$ w.r.t. $(x,\,y)$ coordinates. Again, the changes are from $\mathcal{O}(v_{\infty}^{2})$ onwards and yet again the contributions to $\theta$ and $\rho$ are localised to the boundary of the warp drive.